Fourier Sine Series

Expand F(x)=cosx ; 0<x<π in a Fourier Sine Series. Draw the graph also.
Answer #1

We have to expand the function f(x) = cosx in a Fourier sine series on the interval (OTT). Half range Fourier sine series 0<x2 TT Il-cos(1+1)x (I+h2C So Cos(1-1) X I-n lice jelo bn = 0+ (-COSCI+N)TI+coso) CIENTI )T (-cos(l-n) TT + coso) bn = 0+5)*(1-x2xsin (no) ie f() = E HOT X2 x B1N(136) + (-4)T hoeven noodd (n+1) ( ( By ®, » E 2 sincny) cosx = (1+n) hi even -E 9 0<x<TT